OpenStudy (ny,ny):

Hi =) I am a 7th grader and I'm in an advanced math class at school. 2moro I got a Chapter/Unit test, if I get below an 80 on this one, I will be moved out of the class. SO PLEASE HELP! Thanks y'all. =) **Math SUCKS!!** So is 5x larger than 1/4x by 20...? 0.0. lol help...

hero (hero):

x/4 * 20 = \[\frac{20x}{4} = 5x\]

hero (hero):

5x = 5x

OpenStudy (ny,ny):

so the answer is 20x.

OpenStudy (ny,ny):

cool!

hero (hero):

????

OpenStudy (ny,ny):

...

OpenStudy (ny,ny):

what happen?

hero (hero):

I thought it was obvious, but I guess not

OpenStudy (ny,ny):

what was obvious... the answer? 0.0

hero (hero):

You're in an advanced math class. It should be obvious. But apparently not

OpenStudy (ny,ny):

i did get the answer. I was checking wit u. =(

hero (hero):

The answer is not 20x

OpenStudy (ny,ny):

ok...

OpenStudy (ny,ny):

sorry I was afk.

OpenStudy (ny,ny):

1/4 * 20 = 5...

hero (hero):

What does afk mean?

OpenStudy (ny,ny):

5x is larger than 1/4 by 20

OpenStudy (ny,ny):

away from keyboard

hero (hero):

I think I may have mis-read the question myself. But write it the way you wrote it the first time: 5x is larger than x/4 by 20

OpenStudy (ny,ny):

not x/4 1/4

hero (hero):

x/4 means the same thing as 1/4x. When writing 1/4x linearly, you should write x/4 to avoid confusion.

OpenStudy (ny,ny):

oo

OpenStudy (ny,ny):

thats good to know

OpenStudy (ny,ny):

thank you anyway....